# Anisotropic flow fluctuation as a possible signature of clustered nuclear geometry in O–O collisions at the Large Hadron Collider

### v1

@@ II.1.4 UrQMD: Hadronic cascade
Finally, the produced particles are evolved using the UrQMD microscopic transport model [80, 81]. This model handles both elastic and inelastic processes of hadronic scatterings and resonance decay with particles of masses up to 2.25 GeV. UrQMD simulates a more realistic description of the final-stage hadron evolution and performs a dynamical-freezeout for different particle species. All parameters in the UrQMD model are set to their default values. The output of UrQMD includes final-state particle four momenta and particle identification (PID) code which can be written to disk for further analysis.
II.2 Multiparticle cumulant method
Anisotropic flow can be expressed as the coefficients of Fourier expansion, as shown in Eq. (1). One can estimate the anisotropic flow coefficients using the following equation [82, 83, 84]:
$v_{n}=\langle\cos[n(\phi-\psi_{n})]\rangle\ .$ (3)
@@ II.2 Multiparticle cumulant method
where $m_{p}$ is the total number of particles labeled as POIs, and $m_{q}$ is the total number of particles tagged as both reference flow particles (RFPs) and POIs. RFPs serve as a reference frame for POIs to quantify the collective motion of the system and help establish the orientation of the event plane, which is crucial for determining the anisotropic flow coefficients. The single-event average differential two- and four-particle azimuthal correlation functions are calculated as:
@@
$w_{\langle 4^{{}^{\prime}}\rangle}$ $=(m_{p}M-3m_{q})(M-1)(M-2)(M-3).$
@@
Unfortunately, the $v_{n}$ obtained from the two-particle Q-Cumulant method may contain contributions from non-flow effects, which can be suppressed by appropriate kinematic cuts. One can introduce a pseudorapidity gap between the particles in the two-particle Q-cumulant method [91, 94, 93] to suppress the non-flow contributions. Accordingly, the whole event is divided into two sub-events [95, 96, 97], $A$ and $B$, which are separated by a $|\Delta\eta|$ gap, as depicted in Fig. 3. This modifies Eq. (5) to:
@@
In this paper, the elliptic and triangular flow coefficients are calculated using the above equations by setting $n=$ 2 and 3, respectively. Using Eqs. (9) and (20), one can obtain the mean and fluctuations of anisotropic flow using the following expressions.
$\langle v_{n}\rangle=\sqrt{\frac{v_{n}^{2}\{2\}+v_{n}^{2}\{4\}}{2}}$ (21)
$\sigma_{v_{n}}=\sqrt{\frac{v_{n}^{2}\{2\}-v_{n}^{2}\{4\}}{2}}$ (22)
Using Eqs. (21), and (22), one can obtain relative anisotropic flow fluctuation as follows.
@@
In this study, we use all the charged hadrons within the pseudorapidity region, $|\eta|<2.5$, for the estimation of anisotropic flow coefficients using the two- and four-particle Q-cumulants method. We tag the charged hadrons within $|\eta|<2.5$ and $4.00.2<p_{\rm T}<$ 4.0 GeV/c as RFPs. We apply a pseudorapidity gap, $|\Delta\eta|>1.0$ in the two-subevent method, as shown in Fig. 3, to subtract the non-flow effects from two-particle Q-cumulants method. The statistical uncertainties are calculated using the relations shown in Ref. [98].
@@ III Results and discussions
In this section, we start with a brief discussion of the multiplicity selection, followed by discussions on the results for flow coefficients. Finally, we discuss the centrality dependence of flow fluctuations.
@@
Figure 4: Distribution of charged particle produced in the pseudorapidity acceptance of $-3.7<\eta<-1.7$ and $5.12.8<\eta<$ 5.1 (V0 Multiplicity or V0M) in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model for Woods-Saxon (upper) and $\alpha$-cluster (middle) nuclear density profiles. The lower panel shows the ratio of V0M distributions from Woods-Saxon to that of in $\alpha$-cluster nuclear profile.
@@
Figure 4 shows the distribution of charged particles produced in the pseudorapidity region $-3.7<\eta<-1.7$ and $5.12.8<\eta<$ 5.1 (V0M) in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model for a Woods-Saxon (upper) and $\alpha$-cluster (lower) nuclear density profiles. The lower panel shows the ratio of normalized yields of V0M for Woods-Saxon and $\alpha$-cluster nuclear distributions. Also, different centrality classes are defined based on the percentile slices on the V0M distribution for the Woods-Saxon nuclear density profile. Here, for a fair comparison, we use the same cuts of V0M distribution for all the centrality classes for both the nuclear density profiles. The lowest central class, e.g., (0-10)%, refers to the most central collisions, while the highest centrality class implies a peripheral collision. One can refer to Fig. 8 in the Appendix .1 for the average value of the impact parameter corresponding to each centrality class for both Woods-Saxon and $\alpha$-cluster nuclear density profiles. In Fig. 4, one finds a higher yield for V0M for the O–O collisions having an $\alpha$-clustered nucleus. In addition, for a similar value of V0M, one finds a higher yield for the $\alpha$-clustered distribution on nucleons inside the oxygen nuclei for V0M $>300$, which can be seen in the bottom ratio plot. This signifies the compact nature of an $\alpha$-clustered oxygen nucleus, which is reflected in the most central O–O collisions.
@@
Figure 5: Upper panel shows the elliptic flow and triangular flow using two-particle Q-cumulant method as a function of centrality class (%) in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model for Woods-Saxon and $\alpha$-cluster nuclear density profiles. The lower panel shows the ratio of $v_{n}\{2\}$ for Woods-Saxon to a $\alpha$-cluster nuclear density profile.
@@
Figure 6: Upper panel shows the ratio $v_{3}\{2\}/v_{2}\{2\}$ as a function of centrality class (%) in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model for Woods-Saxon and $\alpha$-cluster nuclear density profiles. Lower panel shows the ratio of $v_{3}\{2\}/v_{2}\{2\}$ from Woods-Saxon to a $\alpha$-cluster nuclear density profile.
@@
Figure 5 shows $v_{2}\{2\}$ and $v_{3}\{2\}$ with a pseudorapidity gap, $|\Delta\eta|>1.0$, as a function of centrality for Woods-Saxon and $\alpha$-cluster nuclear density profiles of 16O in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model. One notices a fair centrality dependence on both elliptic and triangular flow in both Woods-Saxon and $\alpha$-cluster nuclear density profiles. The values of $v_{2}\{2\}$ for the $\alpha$-cluster nuclear density profile are higher than the Woods-Saxon nuclear density profiles throughout the centrality classes except for the peripheral collisions, where the trend is reversed. Here, the value of elliptic flow is highest for the central collisions and decreases towards the peripheral collisions. In most central collisions, the observed high values of elliptic flow can be attributed to the initial eccentricity fluctuations. In contrast, towards the mid-central collisions, although the geometry of the overlap region plays a role, the presence of a lower number of participants makes the initial eccentricity completely transform into the final state elliptic flow of hadrons, thus lowering the value of elliptic flow. However, The observed behavior of elliptic flow as a function of collision centrality and nuclear density profiles assuming a hydrodynamical expansion vary significantly from that observed in a kinetic theory-based model, such as AMPT [6]. In contrast, we observe an interesting behavior for $v_{3}\{2\}$ as a function of collision centrality for both the nuclear density profiles. One finds that $v_{3}\{2\}$ decreases from most central to peripheral collisions, having similar slopes for the Woods-Saxon and $\alpha$-clustered nuclear density profiles within (20-90)centrality classes. This consistency in the triangular flow between two nuclear density profiles for the mentioned centrality class is consistent with the AMPT predictions, as shown in Ref. [6]. Interestingly, it can be observed from the lower panel of Figure 5, a sudden rise of $v_{3}\{2\}$ in the more central collisions for the $\alpha$-clustered in contrast to a Woods-Saxon density profile. These results are similar to the AMPT model calculations in Ref. [6]. We believe that the sudden rise in $v_{3}\{2\}$ for most central collisions of oxygen nuclei having an $\alpha$-clustered nuclear density profile can be attributed to the presence of an initial tetrahedral arrangement of four $\alpha$-particles inside the oxygen nucleus.
@@
Figure 7: Collision centrality dependence of $v_{2}$, $\langle v_{2}\rangle$, $\sigma_{v_{2}}$, and $F(v_{2})$ in O–O collisions, from (a) to (d), respectively, at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model for Woods-Saxon and $\alpha$-cluster nuclear density profiles.
@@
The upper panel of Fig. 6 shows the ratio $v_{3}\{2\}/v_{2}\{2\}$ as a function of collision centrality in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model for Woods-Saxon and $\alpha$-cluster nuclear density profiles. The bottom panel shows the ratio of $v_{3}\{2\}/v_{2}\{2\}$ for Woods-Saxon to a $\alpha$-cluster nuclear density profile. Interestingly, when the oxygen nuclei having the Woods-Saxon nuclear distributions collide, their elliptic and triangular flow values follow a smooth dependence on collision centrality, as one can see in Fig. 5. Thus, the ratio $v_{3}\{2\}/v_{2}\{2\}$ also exhibits a smooth decrease from most central-to-peripheral collisions. However, one finds a notably sharp increase in $v_{3}\{2\}/v_{2}\{2\}$ for (0-5)centrality case for the $\alpha$-cluster case, followed by a smooth decrease in $v_{3}\{2\}/v_{2}\{2\}$ as a function of collision centrality. This reflects the presence of a tetrahedral arrangement of $\alpha$-clusters inside the oxygen nuclei giving rise to a large intrinsic triangular flow in the most central collisions. Although one finds $v_{2}\{2\}$ for $\alpha$-cluster density profile is larger compared to that of Woods-Saxon nuclear density profile, this tetrahedral structure raises $v_{3}\{2\}$ even higher for the most central collisions, which is also consistent with transport models such as AMPT, shown in Ref. [6]. Thus the observation of a sudden rise in tetrahedral $\alpha$-clustered case can lead to an observation of a spike in $v_{3}\{2\}/v_{2}\{2\}$ for most central collisions and consequently, the observation of which in O–O collisions may signify the presence of a $\alpha$-clustered nuclear geometry of 16O nuclei.
@@
Figure 7 (a) shows the evolution of elliptic flow using two- and four-particle Q-cumulant method as a function of centrality for Woods-Saxon and $\alpha$-cluster nuclear density profiles in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model. One finds a small centrality dependence for $v_{2}\{2,|\Delta\eta|>1.0\}$ for both the nuclear density profiles within (0-30)centrality. In contrast, using the four-particle cumulants method, we observe a significant centrality dependence of elliptic flow. The value of $v_{2}\{4\}$ for the $\alpha$-cluster case increases from (0-10)to (10-20)%, and again decreases for (20-30)centrality. However, for the Woods-Saxon case, $v_{2}\{4\}$ maintains a consistent decreasing feature with an increase in the centrality. Another important point to note here is that $v_{2}\{2,|\Delta\eta|>1.0\}>v_{2}\{4\}$ for both the nuclear density profiles, which can consequently lead to real values of flow fluctuations. Figure 7 (b) shows the mean value of elliptic flow ($\langle v_{2}\rangle$), evaluated using Eq. (21), for Woods-Saxon and $\alpha$-cluster nuclear density profiles in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model. $\langle v_{2}\rangle$ is found to be larger for O–O collisions having a $\alpha$-cluster nuclear density profile as compared to a Woods-Saxon distribution. The centrality dependence of $\langle v_{2}\rangle$ for both the nuclear density profiles is similar to $v_{2}\{4\}$, as $v_{2}\{2,|\Delta\eta|>1.0\}$ is nearly independent of collision centrality in (0-30)centrality region. Figure 7 (c) shows a centrality dependence of elliptic flow fluctuation ($\sigma_{v_{2}}$) for Woods-Saxon and $\alpha$-cluster nuclear density profiles in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model. One finds that $\sigma_{v_{2}}$ for Woods-Saxon and $\alpha$-cluster nuclear density profiles have complementing centrality dependence, i.e., when going from most central to mid-central collisions, $\sigma_{v_{2}}$ for Woods-Saxon increases while for an $\alpha$-cluster nuclei it shows a decreasing trend. This makes one of the interesting distinctions between the two nuclear density profiles with the study of flow fluctuations. Although, in Pb–Pb collisions, due to the absence of elliptic geometry in the initial state, the elliptic flow in the most-central collisions has a large contribution from elliptic flow fluctuations; however, towards the mid-central collisions, the contribution of elliptic flow-fluctuations to elliptic flow decreases, as the elliptic geometry of the nuclear participants leads the elliptic flow contribution [93, 96, 97]. Interestingly, the $\alpha$-cluster case of O–O collisions shows a similarity in the evolution of elliptic flow fluctuations with the Pb–Pb collisions, which is clearly different in O–O collisions with a Woods-Saxon nuclear density profile. The source of this opposite behavior of $\sigma_{v_{2}}$ for both the nuclear density profiles can thus intrinsically be attributed to the difference in the nuclear geometry of the 16O nucleus. Interestingly, the centrality dependence of $\langle v_{2}\rangle$ and $\sigma_{v_{2}}$ are complementary to each other for the Woods-Saxon case in O–O collisions. However, this correlation between $\langle v_{2}\rangle$ and $\sigma_{v_{2}}$ is not completely applicable for O–O collisions with an $\alpha$-cluster nuclear density profile. Figure 7 (d) shows the evolution of relative elliptic flow fluctuation ($F(v_{2})$) as a function of collision centrality in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV. Interestingly, we find a similar centrality dependence of $F(v_{2})$ as found for $\sigma_{v_{2}}$. The centrality dependence of $F(v_{2})$ in O–O collisions with Woods-Saxon nuclear distribution resembles a p–Pb collision [93]. In contrast, for $\alpha$-cluster nuclear density profiles, $F(v_{2})$ is higher for most central collisions and decreases towards the mid-central collisions. This is an important observation where we find that a change in the initial nucleon distribution (Woods-Saxon versus $\alpha$-cluster) can cause a difference in the collision centrality dependence of relative elliptic flow fluctuations for a hydro-dynamically evolving system.

### v2

@@ II.1.4 UrQMD: Hadronic cascade
Finally, the produced particles are evolved using the UrQMD microscopic transport model [88, 89]. This model handles both elastic and inelastic processes of hadronic scatterings and resonance decay with particles of masses up to 2.25 GeV. UrQMD closely simulates a realistic description of the final-stage hadron evolution and performs a dynamical freeze-out for different particle species. All parameters in the UrQMD model are set to their default values. The output of UrQMD includes final-state particle four momenta and particle identification (PID) code which can be written to disk for further analysis.
II.2 Multi-particle cumulant method
Anisotropic flow can be expressed as the coefficients of Fourier expansion, as shown in Eq. (1). One can estimate the anisotropic flow coefficients using the following equation [90, 91, 92].
$v_{n}=\langle\cos[n(\phi-\psi_{n})]\rangle$ (3)
@@ II.2 Multi-particle cumulant method
where $m_{p}$ is the total number of particles labeled as POIs, and $m_{q}$ is the total number of particles tagged as both reference flow particles (RFPs) and POIs. RFPs serve as a reference frame for POIs to quantify the collective motion of the system and help establish the orientation of the event plane, which is crucial for determining the anisotropic flow coefficients. The single-event average differential two- and four-particle azimuthal correlation functions are calculated as follows.
@@
$w_{\langle 4^{{}^{\prime}}\rangle}$ $=(m_{p}M-3m_{q})(M-1)(M-2).$
@@
Unfortunately, the $v_{n}$ obtained from the two-particle Q-Cumulant method contains contributions from non-flow effects, which can be suppressed by appropriate kinematic cuts. One can introduce a pseudorapidity gap between the particles in the two-particle Q-cumulant method [96, 95, 97, 99, 98] to suppress the non-flow contributions, which is equivalent to the scalar product method of calculating the anisotropic flow coefficients [100]. Accordingly, the whole event is divided into two sub-events [101, 102, 103], $A$ and $B$, which are separated by a $|\Delta\eta|$ gap, as depicted in Fig. 3. This modifies Eq. (5) to:
@@
In this paper, the elliptic and triangular flow coefficients are calculated using the above equations by setting $n=$ 2 and 3, respectively. Considering small and Gaussian-like fluctuations of anisotropic flow coefficients ($\sigma_{v_{n}}$), the mean ($\langle v_{n}\rangle$) and fluctuations of anisotropic flow coefficients can be represented in terms of $v_{n}\{2,|\Delta\eta|\}$ and $v_{n}\{4\}$, as follows [20, 104].
$\langle v_{n}\rangle=\sqrt{\frac{v_{n}^{2}\{2,|\Delta\eta|\}+v_{n}^{2}\{4\}}{2}}$ (21)
$\sigma_{v_{n}}=\sqrt{\frac{v_{n}^{2}\{2,|\Delta\eta|\}-v_{n}^{2}\{4\}}{2}}$ (22)
Using Eqs. (21), and (22), one can obtain the relative anisotropic flow fluctuation ($F(v_{n})$) as follows.
@@
In this study, we use 3500 IPGlasma + MUSIC events, and each IPGlasma + MUSIC event is sampled 200 times for hadron rescattering using UrQMD. Thus, we have a total of 700K minimum bias events. We use all the charged hadrons within the pseudorapidity region, $|\eta|<0.8$, for the estimation of anisotropic flow coefficients using the two- and four-particle Q-cumulants method. We tag the charged hadrons within $|\eta|<0.8$ and $4.00.2<p_{\rm T}<$ 4.0 GeV/c as RFPs. We apply a pseudorapidity gap, $|\Delta\eta|>1.0$ in the two-subevent method, as shown in Fig. 3, to subtract the non-flow effects from two-particle Q-cumulants method. The statistical uncertainties are calculated using the relations shown in Ref. [105].
@@ III Results and discussions
In this section, we start with a discussion of the multiplicity selection, followed by discussions on the results for anisotropic flow coefficients. Finally, we discuss the multiplicity dependence of flow fluctuations.
@@
Figure 4: Distribution of charged particle produced in the pseudorapidity acceptance of the V0 detector of ALICE, i.e., $-3.7<\eta<-1.7$ and $5.12.8<\eta<$ 5.1 (V0 Multiplicity or V0M) in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model for Woods-Saxon (upper) and $\alpha$-cluster (middle) nuclear density profiles. The lower panel shows the ratio of V0M distributions from Woods-Saxon to that of in $\alpha$-cluster nuclear profile.
@@
Figure 4 shows the distribution of charged particles produced in the pseudorapidity region $-3.7<\eta<-1.7$ and $5.12.8<\eta<$ 5.1 (V0M) in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model for the Woods-Saxon (upper) and $\alpha$-cluster (lower) nuclear density profiles. The lower panel shows the ratio of normalized yields of V0M for Woods-Saxon and $\alpha$-cluster nuclear distributions. Also, different multiplicity classes are defined based on the percentile slices on the V0M distribution for the Woods-Saxon nuclear density profile. Here, for a fair comparison, we use the same cuts of V0M distribution for all the multiplicity classes for both the nuclear density profiles. The first multiplicity class, e.g., 0–10%, has the highest particle multiplicity, while the last multiplicity class, 50–90%, has the events having the lowest final state charged particle multiplicity, as shown in Fig. 4. Thus, one may infer that the 0–10% class refers to the most central collisions, while the last multiplicity class, 50–90% in this study, implies a peripheral collision. Further, one can refer to Fig. 8 in the Appendix .1 for the average value of the impact parameter corresponding to each multiplicity class for both Woods-Saxon and $\alpha$-cluster nuclear density profiles. In Fig. 4, one finds a higher yield for V0M for the O–O collisions having an $\alpha$-clustered nucleus. In addition, for a similar value of V0M, one finds a higher yield for the $\alpha$-clustered distribution of nucleons inside the 16O nuclei for V0M $>300$, which can be seen in the bottom ratio plot. In Ref. [49], the probability distribution of $\alpha$-clustered nuclear density profile of 16O nuclei is compared with Woods-Saxon and Harmonic-Oscillator nuclear density profiles. Here, one finds that the probability of finding nucleons in Harmonic-Oscillator and Woods-Saxon nuclear density profiles spreads over long distances from the center of 16O nuclei. In contrast, the probability of finding nucleons in the $\alpha$-cluster profile is limited to some finite distance from the center of the nucleus. This demonstrates the compactness of the $\alpha$-clustered structure of 16O nuclei, which may lead to increased energy density, as shown in Ref. [49]. A large energy density can lead to an increase in final state particle yield, as can be seen in Fig. 4.
@@
Figure 5: Upper panel shows the elliptic flow and triangular flow using two-particle Q-cumulant method as a function of multiplicity class (%) in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model for Woods-Saxon and $\alpha$-cluster nuclear density profiles. The lower panel shows the ratio of $v_{n}\{2\}$ for Woods-Saxon to a $\alpha$-cluster nuclear density profile.
@@
Figure 6: Upper panel shows the ratio $v_{3}\{2,|\Delta\eta|>1.0\}/v_{2}\{2,|\Delta\eta|>1.0\}$ as a function of multiplicity class (%) in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model for Woods-Saxon and $\alpha$-cluster nuclear density profiles. Lower panel shows the ratio of $v_{3}\{2,|\Delta\eta|>1.0\}/v_{2}\{2,|\Delta\eta|>1.0\}$ from Woods-Saxon to a $\alpha$-cluster nuclear density profile.
@@
Figure 5 shows $v_{2}\{2,|\Delta\eta|>1.0\}$ and $v_{3}\{2,|\Delta\eta|>1.0\}$ with a pseudorapidity gap, $|\Delta\eta|>1.0$, as a function of multiplicity class for Woods-Saxon and $\alpha$-cluster nuclear density profiles of 16O in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model. One notices a moderate multiplicity dependence on both elliptic and triangular flow in both Woods-Saxon and $\alpha$-cluster nuclear density profiles. The values of $v_{2}\{2,|\Delta\eta|>1.0\}$ for the $\alpha$-cluster nuclear density profile are higher than the Woods-Saxon nuclear density profiles throughout the multiplicity classes except for the 50–90% case, where the trend is reversed. This nuclear density profile dependence of $v_{2}\{2,|\Delta\eta|>1.0\}$ has significantly large effects from the initial state eccentricities, as shown in Fig. 9. Further, the value of elliptic flow is the largest for the 0–10% multiplicity class and decreases as the final state particle multiplicity decreases. In the 0–10% multiplicity class, the observed high values of elliptic flow can be attributed to the initial eccentricity fluctuations. In contrast, towards the low multiplicity regions, although the value of eccentricity increases, the continued decrease in the system size and lifetime takes over, restricting the complete transfer of initial eccentricity to the final state azimuthal anisotropy. This results in a smaller value of anisotropic flow coefficients towards the low multiplicity regions than for the high multiplicity events. However, the observed behavior of elliptic flow as a function of multiplicity class and nuclear density profile assuming a hydrodynamical expansion vary significantly from that observed in a kinetic theory-based model, such as AMPT [6]. In contrast, we observe an interesting behavior for $v_{3}\{2,|\Delta\eta|>1.0\}$ as a function of multiplicity class for both the nuclear density profiles. One finds that $v_{3}\{2,|\Delta\eta|>1.0\}$ decreases from 0–10% to 50–90% multiplicity classes, having similar slopes for the Woods-Saxon and $\alpha$-clustered nuclear density profiles within 20–90% multiplicity classes. This consistency in the triangular flow between two nuclear density profiles for the mentioned multiplicity class is consistent with the AMPT predictions, as shown in Ref. [6]. Interestingly, it can be observed from the lower panel of Figure 5, a sudden rise of $v_{3}\{2,|\Delta\eta|>1.0\}$ in the 0–10% multiplicity class for the $\alpha$-clustered in contrast to a Woods-Saxon density profile. These results are similar to the AMPT model calculations in Ref. [6]. We believe that the sudden rise in $v_{3}\{2,|\Delta\eta|>1.0\}$ for 0–10% multiplicity class of 16O nuclei having an $\alpha$-clustered nuclear density profile can be attributed to the presence of an initial tetrahedral arrangement of four $\alpha$-particles inside the 16O nucleus.
@@
Figure 7: Multiplicity class dependence of $v_{2}$, $\langle v_{2}\rangle$, $\sigma_{v_{2}}$, and $F(v_{2})$ in O–O collisions, from (a) to (d), respectively, at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model for Woods-Saxon and $\alpha$-cluster nuclear density profiles.
@@
The upper panel of Fig. 6 shows the ratio $v_{3}\{2,|\Delta\eta|>1.0\}/v_{2}\{2,|\Delta\eta|>1.0\}$ as a function of multiplicity class in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model for Woods-Saxon and $\alpha$-cluster nuclear density profiles. The bottom panel shows the ratio of $v_{3}\{2,|\Delta\eta|>1.0\}/v_{2}\{2,|\Delta\eta|>1.0\}$ for Woods-Saxon to an $\alpha$-cluster nuclear density profile. Interestingly, when the 16O nuclei having the Woods-Saxon nuclear distributions collide, their elliptic and triangular flow values follow a modest dependence on final state multiplicity, as one can see in Fig. 5. Thus, the ratio $v_{3}\{2,|\Delta\eta|>1.0\}/v_{2}\{2,|\Delta\eta|>1.0\}$ also exhibits a modest decrease with a decrease in particle multiplicity. However, one finds a notably stronger increase in $v_{3}\{2,|\Delta\eta|>1.0\}/v_{2}\{2,|\Delta\eta|>1.0\}$ for 0–10% multiplicity class for the $\alpha$-cluster case, followed by a modest decrease in $v_{3}\{2,|\Delta\eta|>1.0\}/v_{2}\{2,|\Delta\eta|>1.0\}$ as a function of multiplicity class. Since these effects originate from the initial state, as shown in Fig. 10, the stronger increase in the value of $v_{3}\{2,|\Delta\eta|>1.0\}/v_{2}\{2,|\Delta\eta|>1.0\}$ in the 0–10% multiplicity class reflects the presence of a tetrahedral arrangement of four $\alpha$-clusters inside the 16O nuclei, giving rise to a large intrinsic triangular flow. Although one finds $v_{2}\{2,|\Delta\eta|>1.0\}$ for $\alpha$-cluster density profile is larger compared to that of Woods-Saxon nuclear density profile, this tetrahedral structure raises $v_{3}\{2,|\Delta\eta|>1.0\}$ even higher for the high-multiplicity events, which is also consistent with transport models such as AMPT, shown in Ref. [6]. Thus, the observation of a sudden rise of $v_{3}\{2,|\Delta\eta|>1.0\}$ in tetrahedral $\alpha$-clustered case can lead to an observation of a spike in $v_{3}\{2,|\Delta\eta|>1.0\}/v_{2}\{2,|\Delta\eta|>1.0\}$ for the highest charged particle multiplicity events and consequently, the observation of which in O–O collisions may signify the presence of a $\alpha$-clustered nuclear geometry of 16O nuclei.
@@
Figure 7 (a) shows the evolution of elliptic flow using two- and four-particle Q-cumulant method as a function of multiplicity class for Woods-Saxon and $\alpha$-cluster nuclear density profiles in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model. One finds a small multiplicity dependence for $v_{2}\{2,|\Delta\eta|>1.0\}$ for both the nuclear density profiles within 0–30% multiplicity class. In contrast, using the four-particle cumulants method, we observe a significant charged particle multiplicity dependence of elliptic flow. The value of $v_{2}\{4\}$ for the $\alpha$-cluster case remains almost unchanged as a function of multiplicity class, the same trend as of $v_{2}\{2\}$. However, for the Woods-Saxon case, $v_{2}\{4\}$ maintains a consistent decreasing feature with the decrease in the particle multiplicity. Figure 7 (b) shows the mean value of elliptic flow ($\langle v_{2}\rangle$), evaluated using Eq. (21), for Woods-Saxon and $\alpha$-cluster nuclear density profiles in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model. $\langle v_{2}\rangle$ is found to be larger for O–O collisions having an $\alpha$-cluster nuclear density profile as compared to a Woods-Saxon distribution. The multiplicity class dependence of $\langle v_{2}\rangle$ for both the nuclear density profiles is similar to $v_{2}\{4\}$, as $v_{2}\{2,|\Delta\eta|>1.0\}$ is nearly independent of multiplicity class in the 0–30% region. Figure 7 (c) shows a multiplicity class dependence of elliptic flow fluctuation ($\sigma_{v_{2}}$) for Woods-Saxon and $\alpha$-cluster nuclear density profiles in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model. One finds that $\sigma_{v_{2}}$ for Woods-Saxon and $\alpha$-cluster nuclear density profiles have complementing multiplicity dependence, i.e., when going from the 0–10% to 20–30% multiplicity classes, $\sigma_{v_{2}}$ for Woods-Saxon increases while for an $\alpha$-cluster nuclei, it shows a decreasing trend. This makes one of the interesting distinctions between the two nuclear density profiles with the study of flow fluctuations. In Pb–Pb collisions, due to the absence of the elliptic geometry in the initial state, a larger contribution to the elliptic flow in the most central collisions comes from elliptic flow fluctuations; however, towards the mid-central collisions, the contribution of elliptic flow-fluctuations to elliptic flow decreases, as the elliptic geometry of the nuclear participants leads the elliptic flow contribution [98, 102, 103]. Interestingly, the $\alpha$-cluster case of O–O collisions shows a similarity in the evolution of elliptic flow fluctuations with the Pb–Pb collisions, which is clearly different in O–O collisions with a Woods-Saxon nuclear density profile. The source of this opposite behavior of $\sigma_{v_{2}}$ for both the nuclear density profiles can thus intrinsically be attributed to the difference in the nuclear geometry of the 16O nucleus. Interestingly, the multiplicity class dependence of $\langle v_{2}\rangle$ and $\sigma_{v_{2}}$ are complementary to each other for the Woods-Saxon case in O–O collisions. However, this correlation between $\langle v_{2}\rangle$ and $\sigma_{v_{2}}$ is not completely applicable for O–O collisions with an $\alpha$-cluster nuclear density profile. Figure 7 (d) shows the evolution of relative elliptic flow fluctuation ($F(v_{2})$) as a function of multiplicity class in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV. Interestingly, we find a similar multiplicity dependence of $F(v_{2})$ as found for $\sigma_{v_{2}}$. The multiplicity class dependence of $F(v_{2})$ in O–O collisions with Woods-Saxon nuclear distribution resembles a p–Pb collision [98]. In contrast, for $\alpha$-cluster nuclear density profiles, $F(v_{2})$ is higher for 0–10% multiplicity class and decreases with a decrease in particle multiplicity. This is an important observation where we find that a change in the initial nucleon distribution (Woods-Saxon versus $\alpha$-cluster) can cause a difference in the multiplicity class or collision centrality dependence of relative elliptic flow fluctuations for a hydrodynamically evolving system.